# Make Autoregressive Great Again: Diffusion-Free Graph Generation with Next-Scale Prediction

### v1

@@ Discussion
Multi-scale prediction offers substantial properties that allow MAG to address the three previously mentioned issues as follows:
@@
    Given the limited number of intermediate scales in the generations, long-range dependency is reduced from $\mathcal{O}(n^{2})$ to $\mathcal{O}(k)$, where $k\approx\log(n)$ is the number of scales. Furthermore, sequential dependency is strengthened by generating a scale of the whole graph at each autoregressive step.
@@
    No explicit node ordering is required, as (i) the tokenizer is permutation equivariant, and (ii) tokens in each map $r_{k}$ are fully correlated and generated in parallel.
@@
    By generating an entire token map in parallel at each scale, the generation’s complexity is reduced from $\mathcal{O}(n^{3})$ to $\mathcal{O}(kn^{2})$, where $k\approx\log(n)$. The proof is detailed in Appendix A.
@@ 4 Experiments
We empirically evaluate our model on experiments covering both generic and molecular graph generation. These two domains are chosen to benchmark the model’s versatility: generic graphs allow the assessment of structural fidelity and scalability, while molecular graphs pose domain-specific challenges like chemical validity and functional constraints. This dual setting probes MAG’s ability to generate diverse, high-quality graphs under different structural and semantic constraints.
Generation quality is measured using a combination of distributional and domain-specific metrics. Moreover, we evaluate generation time against state-of-the-art diffusion, hybrid, and VAE models to compare their inference efficiency.
@@
We compare MAG to state-of-the-art models trained for both generic and molecule generation. Specifically, baselines include diffusion models, such as EDP-GNN [6], GDSS [7], and DiGress [8]; hybrid models, such as GraphAF [15], GraphDF [24] and GraphRNN [3] (generic graphs only); and VAE models, such as GraphVAE [25] and GramVAE [26] (molecular graphs only).
Figure 4: Samples of generated generic and molecular graphs by MAG.
@@ 4.1 Generic Graph Generation
We evaluate MAG at generic graph generation on the Community-small [7] and Enzymes [27] datasets. The former is a synthetic dataset containing 100 random graphs consisting of two equal-sized communities, with between 12 and 20 nodes per graph; while the latter contains 587 protein graphs with 10 to 125 nodes per protein.
These datasets provide a wide range of graph properties, from small structured communities in Community-small to larger and more complex protein graphs in Enzymes. We aim to investigate the ability of multi-scale graph generation to generate graphs of different sizes and complexities, providing insights into its scalability and generalization abilities.
To ensure fair comparison, we follow the same setup as [9, 11, 7], randomly splitting the dataset into 80% for training and 20% for testing. Performance is evaluated by comparing the structure of the generated graphs with those from the dataset according to the maximum mean discrepancy (MMD) of statistics including node degrees, clustering coefficients, and orbit counts. Furthermore, the inference speed of MAG is compared to the baselines at generating graphs of varying sizes. Samples of generated graphs are displayed in Figure 4.
Results
Table 1 presents the quantitative MMD results comparing our proposed method to baseline models. MAG demonstrates the ability to generate high-quality samples that closely resemble the original data. On the smaller Community-small dataset, MAG consistently outperforms hybrid models and VAEs. Notably, its performance approaches that of state-of-the-art GDSS and DiGress models.
For the more complex Enzymes dataset, MAG further surpasses hybrid and VAE baselines in both performance and computational efficiency. Hybrid approaches struggle with larger enzymes, highlighting scalability limitations, whereas MAG enjoys strong scalability from its autoregressive nature. Consequently, our method achieves performance comparable to diffusion-based approaches on these.
In terms of computational efficiency, inference time evaluation reveals significant improvements across all datasets, surpassing all baselines by several orders of magnitude. Remarkably, compared to state-of-the-art models, MAG achieves inference speeds of up to thousands of times faster.
Table 1: Quantitative results on Community-small and Enzymes benchmark datasets. Inference times are reported in seconds for 50 and 100 samples, respectively, on an NVIDIA GeForce RTX 4060 GPU. Orbit results are missing due to computational incompatibility.
@@ Results
Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Time $\downarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Time $\downarrow$
@@
Ours 0.077 0.102 0.187 1.753 1.148 1.88
@@ 4.2 Molecule Generation
MAG is further extended to generate molecular graphs with node and edge features, and evaluated on two popular molecular datasets, namely QM9 [28] and ZINC-250K [29]. The former contains 133,885 small organic molecules of up to nine heavy atoms with types C, O, N and F. The latter is a subset of the ZINC database and consists of 250,000 molecules with up to 38 heavy atoms of nine types.
@@
The quality of generated molecules is measured in terms of validity, uniqueness, novelty, and Fréchet ChemNet Distance (FCD). Details on these metrics are delineated in Appendix B. In addition, the inference time to generate 1,000 molecules from both datasets is recorded and compared against baselines. Samples of generated molecules are displayed in Figure 4.
@@ Results
Table 2: Comparison of generation times using NVIDIA GeForce RTX 4060 on molecular datasets. Times are reported in seconds for 1,000 samples.
Models Molecular graphs
QM9 ZINC250K
EDP-GNN $1.74e^{3}$ $1.62e^{3}$
GDSS $43.4$ $3.44e^{2}$
DiGress $34.1$ $2.03e^{2}$
GraphAF $1.05e^{3}$ $9.61e^{2}$
GraphDF $1.93e^{4}$ $9.29e^{3}$
Ours 3.84 19.7
Qualitative results on molecular datasets are given in Table 3. These results showcase MAG’s ability to generate high-quality molecules that respect imposed chemical constraints. As observed, the proposed method surpasses VAE baselines across most metrics. Furthermore, MAG enjoys performance competitive to those of hybrid models.
As for the larger, more complex ZINC-250k dataset, MAG surpasses VAE baselines across all metrics and offers comparable results to hybrid models, falling just short of those posed by diffusion models. These results demonstrate that, although MAG is able to effectively generate realistic molecules while adhering to chemical structural constraints, its performance is slightly reduced on complex molecules.
With regards to computational efficiency, drastic improvements in inference time evaluation remain consistent for molecular graphs, surpassing all baselines by at least one order of magnitude and achieving inference speeds of up to hundreds of times faster.
Table 3: Quantitative results on QM9 and ZINC250k molecule datasets.
@@
Valid $\uparrow$ Unique $\uparrow$ Novel $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ Novel $\uparrow$ FCD $\downarrow$
@@
Ours 80.28 99.68 83.14 3.42 77.27 86.27 100 16.88
5 Future Work
This work primarily focused on introducing a novel paradigm for autoregressive graph generation, adopting a next-scale prediction approach instead of traditional next-node and next-edge methods. As a result, a promising avenue for future research is to explore how autoregressive properties from the large language model (LLM) literature translate to graph generation. Specifically, efforts could be made to investigate the generalization capabilities of autoregressive models in zero-shot and few-shot downstream tasks, as well as their scalability to larger graphs.
Additionally, incorporating inductive biases into the graphs’ scales based on domain-specific knowledge or additional contextual information could further enhance model performance and applicability.
Finally, this work opens a new direction for hybrid autoregressive-diffusion methods. While current hybrid approaches primarily rely on block-wise diffusion with sequential learning between blocks, the proposed framework enables reconsidering these methods by applying scale-wise diffusion, i.e., diffusion between scales. This shift could eliminate the need for quantization, as diffusion offers strong capabilities in continuous space, thereby removing the autoencoder’s quantization bottleneck. However, this would reintroduce computational overheads which should be studied against existing diffusion-based approaches.
@@ 6 Conclusion
We introduced a novel diffusion-free multi-scale autoregressive model for graph generation which (1) theoretically addresses limitations of traditional autoregressive (AR) models due to inherent graph properties, and (2) reduces the gap between diffusion-free transformer-based AR models and diffusion-based methods, with up to three orders of magnitude lower computational costs. Experiments over both generic and molecular graph generations demonstrate the ability of MAG to generate high-quality samples that achieve competitive performance with state-of-the-art models and adjust to domain-specific constraints such as chemical validity. We hope that our findings can aid the design of new approaches, in which the multi-scale paradigm may be applied to other graph generative methods.
Acknowledgements
We would like to thank Miruna Cretu, Charlie Harris, and Karolis Martinkus for their guidance and valuable feedback throughout this work.
Appendix A Proof of Complexity Reduction
@@
Lemma 1 (AR Generation for Nodes).
@@
Lemma 2 (Scale-wise Generation for Nodes).
@@
Appendix B Details for Evaluation Metrics
@@
We evaluate the performance on molecular graphs by measuring uniqueness, novelty, validity, FCD, and NSPDK. Uniqueness and novelty ensure that the model can extrapolate to new molecules rather than overfitting the training data, while validity verifies whether generated molecules follow fundamental chemical rules, such as valency constraints. In addition, FCD measures how closely the generated molecules resemble real compounds in terms of learned chemical properties, while NSPDK evaluates structural similarity at the subgraph level.

### v2

@@ Discussion
As a result, multi-scale prediction offers substantial properties that allow MAG to address the three previously mentioned issues as follows:
@@
    Given the limited number of intermediate scales in the generations, long-range dependency is reduced from $\mathcal{O}(n)$ to $\mathcal{O}(k)$, where $k\approx\log(n)$ is the number of scales. Furthermore, sequential dependency is strengthened by generating a scale of the whole graph at each autoregressive step.
@@
    No explicit node ordering is needed as (i) the tokenizer is permutation equivariant, and (ii) tokens in each map, $r_{k}$, are fully correlated and generated in parallel.
@@
    By generating an entire token map in parallel at each scale, the generation’s complexity is reduced from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$. The proof is detailed in Appendix B.
@@ 4 Experiments
We empirically evaluate our model on experiments covering both generic and molecular graph generation. These two domains are chosen to benchmark the model’s versatility: generic graphs allow the assessment of structural fidelity and scalability, while molecular graphs pose domain-specific challenges like chemical validity and functional constraints. This dual setting probes MAG’s ability to generate diverse graphs under different structural and semantic constraints.
Generation quality is measured using a combination of distributional and domain-specific metrics. Moreover, we evaluate generation wall-clock time against state-of-the-art diffusion, hybrid, and VAE models to compare their inference efficiency. Both training and evaluation procedures were executed in five independent trials, with error bars denoting the robustness of the experiments.
@@
We compare MAG to state-of-the-art models trained for both generic and molecule generation. Specifically, baselines include diffusion models, such as EDP-GNN (Niu et al., 2020), GDSS (Jo et al., 2022), and DiGress (Vignac et al., 2023); hybrid models, such as GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021) and GraphRNN (You et al., 2018) (generic graphs only); and VAE models, such as GraphVAE (Simonovsky & Komodakis, 2018) and GramVAE (Kusner et al., 2017) (molecular graphs only).
@@ 4.1 Generic Graph Generation
We evaluate MAG at generic graph generation on the Community-small (Jo et al., 2022) and Enzymes (Schomburg et al., 2004) datasets. Community-small is a synthetic dataset containing 100 random graphs consisting of two equal-sized communities, with between 12 and 20 nodes per graph. Here, the node and edge representations are binary (i.e., absence or presence). The Enzymes dataset contains 587 enzyme graphs with 10 to 125 nodes per sample. Each of the graphs represents a protein as a set of nodes—each node carrying a 3-dimensional one-hot feature for $\alpha$-helix, $\beta$-strand or coil—and binary edges indicating sequence adjacency or the three nearest spatial neighbors. There are no edge attributes, and each graph is annotated with one of six top-level EC enzyme class labels.
These datasets provide a wide range of graph properties, from small structured communities in Community-small to larger, more complex protein graphs in Enzymes. We aim to investigate the ability of multi-scale graph generation to generate graphs of different sizes and complexities, providing insights into its scalability and generalization abilities.
To ensure fair comparison, we follow the same setup as (Kong et al., 2023; Yan et al., 2024; Jo et al., 2022), randomly splitting the dataset into 80% for training and 20% for testing. Performance is evaluated by comparing the structure of the generated graphs with those from the dataset according to the maximum mean discrepancy (MMD) of statistics including node degrees, clustering coefficients, and orbit counts. Moreover, the inference speed of MAG is compared to the baselines at generating graphs across varying sizes. Samples of generated graphs are given in Figure 4.
Table 1: Results on Community-small and Enzymes benchmark datasets. Wall-clock inference times are reported in seconds for 50 and 100 samples, respectively, on an NVIDIA GeForce RTX 4060 GPU.
@@
Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Time (s) $\downarrow$ Deg. $\downarrow$ Clus. $\downarrow$ Orbit. $\downarrow$ Time (s) $\downarrow$
@@
Ours 0.077 0.102 0.022 0.187 1.753 1.148 0.223 1.88
$\pm$ std. $\pm$0.007 $\pm$0.009 $\pm$0.003 $\pm$0.015 $\pm$0.150 $\pm$0.100 $\pm$0.034 $\pm$0.200
Results
[…]
On the more complex Enzymes dataset, MAG exhibits limitations in qualitative performance. Similarly to hybrid models, it struggles with generating larger enzyme graphs, suggesting possible scalability challenges. This may be due to the increased number of scales required for larger graphs, resulting in higher sensitivity to predefined scale sizes. Despite this limitation, MAG’s computational efficiency scales robustly with graph size, indicating strong potential should its performance be further improved.
In terms of wall-clock inference time, MAG delivers significant improvements across all datasets, surpassing all baselines by several orders of magnitude. Remarkably, it achieves inference speeds up to thousands of times faster than state-of-the-art models.
@@ 4.2 Molecule Generation
MAG is further extended to generate molecular graphs with node and edge features, and evaluated on two popular molecular datasets, namely QM9 (Ramakrishnan et al., 2014) and ZINC-250K (Gómez-Bombarelli et al., 2018). The former contains 133,885 small organic molecules of up to nine heavy atoms with types C, O, N and F. The latter is a subset of the ZINC database (Irwin et al., 2012) and consists of 250,000 molecules with up to 38 heavy atoms of nine types.
@@
The quality of generated molecules is measured in terms of validity, uniqueness, novelty, and Fréchet ChemNet Distance (FCD). Details on these metrics are delineated in Appendix C. Training and evaluation were each performed in five independent replicates, with the resulting performance robustness illustrated by error bars. In addition, the wall-clock inference time to generate 1,000 molecules from both datasets is recorded and compared against baselines. Samples of generated molecules are displayed in Figure 4.
@@ Results
Performance on molecular datasets are given in Table 2. MAG consistently outperforms VAE baselines across most metrics. Moreover, MAG achieves performance comparable to that of recent hybrid models. Compared to state-of-the-art diffusion models, MAG achieves the highest uniqueness and matches most of them in novelty. Regarding validity, although it trails slightly behind GDSS, it proves to outperform EDP-GNN. Of all baselines, only DiGress exhibits a notable performance advantage over MAG; however, DiGress explicitly incorporates molecular-specific properties, while MAG approaches the task in a raw, domain-agnostic manner, which may account for this performance gap. For the larger and more complex ZINC-250k dataset, MAG outperforms VAE baselines across all metrics and delivers results comparable to those of hybrid models, while remaining slightly behind diffusion-based methods. These findings are consistent with observations on smaller molecules; however, MAG demonstrates greater consistency across different molecule sizes compared to diffusion models. Additionally, we observe reduced overhead when scaling to larger molecules relative to the results on the Enzymes dataset. Regarding computational efficiency, MAG further delivers substantial improvements, achieving inference speeds up to hundreds of times faster than the state of the art.
Table 2: Results on QM9 and ZINC250k molecule datasets. Wall-clock inference times are reported in seconds for 1,000 samples, respectively, on an NVIDIA GeForce RTX 4060 GPU (Val: Validity; Uni: Uniqueness; Nov: Novelty; FCD: Fréchet ChemNet Distance). Entries marked “–” denote results that were not reported or for which reproduction was not possible.
@@
Val. $\uparrow$ Uni. $\uparrow$ Nov. $\uparrow$ FCD $\downarrow$ Time (s) $\downarrow$ Val. $\uparrow$ Uni. $\uparrow$ Nov. $\uparrow$ FCD $\downarrow$ Time (s) $\downarrow$
@@
Ours 89.43 99.31 83.14 5.77 3.84 84.22 87.88 100 19.70 19.7
$\pm$ std. $\pm$0.87 $\pm$0.38 $\pm$1.72 $\pm$0.49 $\pm$1.40 $\pm$1.32 $\pm$0.53 $\pm$0.00 $\pm$1.06 $\pm$0.58
5 Future Work and Limitations
This work primarily focused on introducing a novel paradigm for autoregressive graph generation, adopting a next-scale prediction approach instead of traditional next-node and next-edge methods. As a result, a promising avenue for future research is to explore how autoregressive properties from the large language model (LLM) literature translate to graph generation. Specifically, efforts could be made to investigate the generalization capabilities of autoregressive models in zero-shot and few-shot downstream tasks, which have been shown to be limitations of state-of-the-art diffusion-based graph models (Vignac et al., 2023). Furthermore, future work could explore fine-tuning strategies as a solution to tasks with little data and resources.
Although MAG demonstrated substantial speed improvements, experiments on larger graphs revealed qualitative limitations—likely stemming from our use of a fixed set of scales. Incorporating inductive biases or domain-specific signals into scale selection could enhance both performance and scalability on more complex structures. Moreover, autoregressive models are known for their inherent scalability, suggesting that MAG’s performance on larger graphs could further improve through optimized scale selection. A promising direction is to train a unified decoder that maps any intermediate latent scale back to a graph, much like Jiao et al.’s multi-resolution VAE. This would not only allow interpretation and visualization of the coarse-to-fine hierarchy and how motifs emerge, but also drive a data-driven choice for scale counts and resolutions, alleviating our current dependence on manually tuned scale hyperparameters.
Finally, this work opens a new direction for hybrid autoregressive-diffusion methods. While current hybrid approaches primarily rely on block-wise diffusion with sequential learning between blocks, the proposed framework enables reconsidering these methods by applying scale-wise diffusion, i.e., diffusion between scales. This shift could eliminate the need for a quantized autoencoder, as diffusion offers strong capabilities in continuous space, thereby removing the quantization bottleneck. However, this would reintroduce computational overheads which should be studied as a trade-off against existing diffusion-based methods.
@@ 6 Conclusion
This work introduced a novel diffusion-free multi-scale autoregressive model for graph generation which (1) theoretically addresses limitations of traditional autoregressive models due to inherent graph properties and (2) reduces the gap between diffusion-free AR models and diffusion-based methods, with up to three orders of magnitude lower computational costs. Experiments over both generic and molecular graph generations demonstrated the ability of MAG to generate high-quality samples that achieve promising performance and can adjust to domain-specific constraints such as chemical validity. We hope that our findings can aid the design of new approaches, in which the multi-scale paradigm may be applied to other graph generative methods.
Impact Statement
Fast graph generation can accelerate scientific discovery by enabling rapid simulation of complex networks in fields such as epidemiology, ecology, and social science, lowering barriers for smaller research teams. It supports more resilient infrastructure and public‐health planning by allowing practitioners to model extreme scenarios at unprecedented speed. By democratizing access to advanced network modeling, these tools empower communities worldwide to tackle interconnected societal challenges.
Acknowledgment
We thank Martinkus Karolis for his detailed feedback on architectural design, his comments on final drafts, and the provision of key literature references that strengthened our work. We also extend our gratitude to Petar Veličković for his early feedback on the foundations of this work and his continued support throughout.
Appendix A Details for Hyperparameters
Table 3: Hyperparameters used for MAG during training and inference.
Hyperparameter Value
Multi‐Scale VQ‐VAE Tokenizer
Encoder MPNN layers ($L_{e}$) 4
Decoder GCN layers ($L_{d}$) 4
Hidden dimension 32
Latent dimension ($C$) 16
Codebook size ($V$) 1024
Commitment cost 0.25
Gamma 0.1
Multi-scale Transformer
Transformer blocks 8
Hidden size 256
Attention heads 8
Level embedding dim 256
Layer dropout 0.1
Conditional dropout 0.1
Token dropout 0.05
Optimization & Sampling
Optimizer Adam
Learning rate $3\times 10^{-5}$
Weight decay $1\times 10^{-2}$
Betas (0.9, 0.99)
Batch size 12
Training epochs 100
Top-$k$ sampling ($k$) 50
Top-$p$ sampling ($p$) 0.95
Appendix B Complexity Proof
@@
Lemma B.1 (AR Generation for Nodes).
@@
Lemma B.2 (Scale-wise Generation for Nodes).
@@
Appendix C Details for Evaluation Metrics
@@
We assess molecular graph generation by examining validity, uniqueness, novelty, and the Frechet ChemNet Distance (FCD) metric. Validity reflects the proportion of generated structures that adhere to fundamental chemical rules—such as correct valency—without any post-hoc correction. Uniqueness gauges the degree to which those valid molecules are distinct from one another, while novelty captures the extent to which they lie outside the original training corpus. Finally, FCD quantifies how closely the distribution of generated compounds matches that of real chemicals by comparing the mean and covariance of their activations in ChemNet’s penultimate layer.